# Local EGOP for Continuous Index Learning

### v2

@@ 3 Bias Control via EGOP
In this section, we develop a theoretical framework for adaptive kernel machines. Define the empirical kernel estimator for predicting the value of $f$ at $x^{*}$ by
$\hat{P}_{M}(f)$ $:=\frac{1}{\hat{C}_{M}}\frac{1}{n}\sum_{i=1}^{n}k(\|M^{1/2}[X_{i}-x^{*}]\|/\sqrt{2})Y_{i},$ (1)
$\hat{C}_{M}$ $:=\frac{1}{n}\sum_{i=1}^{n}k(\|M^{1/2}[X_{i}-x^{*}]\|/\sqrt{2}),$ (2)
where the $D\times D$ matrix $M\succeq 0$ is a Mahalanobis metric to be learned. Its continuous counterpart is
$P_{M}(f)$ $:=\frac{1}{C_{M}}\int k(\|M^{1/2}[y-x^{*}]\|/\sqrt{2})f(y)\,dP(y),$ (3)
$C_{M}$ $:=\int k(\|M^{1/2}[y-x^{*}]\|/\sqrt{2})\,dP(y).$ (4)
@@
To assess the quality of this predictor, we can measure its bias $\int(P_{M}(f)-f)^{2}\,d\mu$ on a target distribution $\mu$. Our key insight is that, when $M$ and $\mu$ are chosen in correspondence, the bias can be related to a Dirichlet form in terms of the EGOP. To make this explicit, we introduce the EGOP-form
@@
Let $M_{t}\succeq 0$, and set $\mu_{t}=N(0,M_{t}^{-1})$. Then,
$\int(P_{M_{t}}(f)-f)^{2}\,d\mu_{t}=O(W(\mu_{t})).$
Thus, if the anisotropic metric $M_{t}$ is such that the variation of $f$ around $x^{*}$, given by $\mathcal{L}(\mu_{t})$, and the variance of $x$ itself, given by $\Sigma(\mu_{t})$, are close to low-rank and orthogonal, then the bias of the kernel estimator (3) is controlled by $W(\mu_{t})=\langle\mathcal{L}(\mu_{t}),\Sigma(\mu_{t})\rangle_{F}$, and will be nearly 0. Of critical importance are the first and second moments of the target distribution $\mu$, motivating our Gaussian convention, although Gaussianity is not essential to achieve the desired bound. We note that this general structure, of metrizing by the inverse covariance, appears in disparate literatures, from classical local feature estimators (Schmid and Mohr, 1997), to VAEs (Chadebec and Allassonnière, 2022), and Langevin preconditioning (Cui et al., 2024).
@@
Let $M_{t}\succeq 0$ and set $\Sigma_{t}=M_{t}^{-1}$, $\mu_{t}=N(x^{*},\Sigma_{t})$, . Then $\mathbb{E}\left[(\hat{P}_{M_{t}}(f)-P_{M_{t}}(f))^{2}\right]=O\left(1/(n\sqrt{\det\Sigma_{t}})\right).$
@@
Let $M_{t}\succeq 0$, and set $\Sigma_{t}=M_{t}^{-1}$, $\mu_{t}=N(x^{*},\Sigma_{t})$. Then,
$\mathbb{E}\left[\int(\hat{P}_{M_{t}}(f)-f)^{2}\,d\mu_{t}\right]=O\left(\frac{1}{n\sqrt{\det\Sigma_{t}}}+W(\mu_{t})\right).$
In particular, if $W(\mu_{t})=O(t)$ and $\det\Sigma_{t}=O(t^{q})$ then the asymptotic learning rate is
$\mathbb{E}\left[\int(\hat{P}_{M_{t}}(f)-f)^{2}\,d\mu_{t}\right]=O\left(n^{-\frac{1}{1+q/2}}\right).$
Thus, if we design a sequence $\{t_{i}\}_{i\in\mathbb{N}}$, with $t_{i}\to 0$, such that $M_{t_{i}}$ and $\mu_{t_{i}}$ are updated recursively based on their values up to $t_{i-1}$ and $\frac{1}{n\sqrt{\det\Sigma_{t_{i_{n}}}}}+W(\mu_{t_{i_{n}}})\to 0$, the kernel estimator $\hat{P}_{M_{t_{i_{n}}}}$ is consistent and its rate is controlled by the convergence rate of $W$ and $M$.
4 The Local EGOP Algorithm
In this section, we introduce the Local EGOP Learning algorithm, detailed in Algorithm 1. The theoretical motivation for this iterative scheme is presented in Section 5. The algorithm iteratively metrizes a kernel by an AGOP, leveraging this operator to assign weights to points neighboring $x^{*}$.
Input: Data $\{(X_{i},Y_{i})\}_{i=1}^{n}$, target $x^{*}$, subsample size $m$, iterations $T$, bandwidths $\{t_{i}\}$, $\beta>0$, scale $\alpha>0$.
Output: Best estimate $\widehat{f}(x^{*})$.
Hyperparameters: $M_{0}\leftarrow I/\alpha$.
1
[…]
3
4for $i=0$ to T1T-1 do
 $\triangleright$ Compute global weights at $x^{*}$
5 $w_{j}\propto\exp\!\bigl(-(X_{j}-x^{*})^{\top}M_{i}(X_{j}-x^{*})\bigr),\ j=1,\dots,n$;
6   Normalize $w$, subsample $S\subset\{1,\dots,n\},|S|=m$ with probs $w$;
7
 $\triangleright$ Leave-One-Out Local regressions at each $j\in S$
8 for $j\in S$, do
9      Fit weighted local linear regression centered at $X_{j}$ (weights from metric $M_{i}$), data point $(X_{j},Y_{j})$ excluded;
10      Obtain $\widehat{\nabla}f[j]$ and $\widehat{f}[j]$;
11
12
 $\triangleright$ Aggregate statistics
13 $L_{i}\leftarrow\sum_{j\in S}w_{j}\,\widehat{\nabla}f[j]\widehat{\nabla}f[j]^{\top}$;
14 $\widehat{\mathrm{MSE}}_{i}\leftarrow\sum_{j\in S}w_{j}\,(Y_{j}-\widehat{f}[j])^{2}$;
15
16 if $\widehat{\mathrm{MSE}}_{i}<\text{bestMSE}$ then
17      Estimate $\widehat{f}(x^{*})$ via local regression centered at $x^{*}$;
18    $\text{bestMSE}\leftarrow\widehat{\mathrm{MSE}}_{i}$,  $\widehat{f}_{\text{best}}(x^{*})\leftarrow\widehat{f}(x^{*})$;
19
20
 $\triangleright$ Metric update: set $\beta=1$ if $i=0$, $\beta<1$ otherwise
21 $M_{i+1}\leftarrow\bigl(\beta L_{i}+(1-\beta)L_{i-1}\bigr)$;
22 $M_{i+1}\leftarrow M_{i+1}/[t_{i+1}\operatorname{tr}(M_{i+1})]$;
23
24
return $\widehat{f}_{\text{best}}(x)$
Algorithm 1 Local EGOP Learning
The time complexity of the method is $O(T((m+2)nD^{2}+(m+1)D^{3}))$, with memory overhead $O(nD^{2}+m^{2}D^{2}+mD^{3})$.
Incorporated into this algorithm is a one-step lookback (step 13) facilitated through $\beta\in(0,1)$. This is essential to improve stability, as discussed in Example 2. In practice, one might choose hyper-parameters $\beta,T$, as well as the schedule $\{t_{i}\}$ via cross-validation. Additionally, for computational reasons, we normalize $M_{i}$ by its trace (step 14) before applying the scaling factor. This has little impact asymptotically, as the trace converges to the leading gradient eigenvalue of the EGOP, making this normalization effectively equivalent to normalization by a constant scalar. However, it serves a practical benefit by stabilizing initial iterates, setting the scale of the metric to the predefined rate $1/t_{i}$. Note that while the optimal number of iterations is dimension dependent, as argued in Theorems 1 and 3, our algorithm incorporates a Leave-One-Out style validation loop (step 10) for adaptive stopping.
5 Recursive Kernel Learning as Variance Minimization
@@
We will now show how recursive kernel learning, combined with our basic kernel bounds, constitutes a greedy variance reduction strategy. Intuitively, the goal is to induce anisotropy in the kernel, stretching or steering it to include additional low-bias points that would otherwise be discarded with a RBF. The inclusion of these additional high-quality points is exactly a reduction of variance, as including more data results in tighter concentration to the mean for a local averaging estimator.
@@
We now make precise how one can arrive at this by leveraging Theorem 1. Ideally, as spelled out in this result, we would like to select $\mu_{t}$ such that $W(\mu_{t})\to 0$, and that $\det\Sigma(\mu_{t})$ is as large as possible. In other words, if we specify the rate $W(\mu_{t})=\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]=t\to 0$, it would be optimal to select the largest possible region that achieves this bias, meaning we would like to select
$\mu_{t}=\operatorname{argmax}_{\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]\leq t}\log\det\Sigma(\mu_{t}).$
@@
This, however, cannot be solved, as we lack precise knowledge of the function of interest $f$. Moreover, $\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]\leq t$ presents a challenging feasible set, as it is posed over a space of probability distributions, thus making it non-Euclidean. Our Gaussian ansatz reduces the problem to optimizing a finite parameterization, however this still presents a non-convex optimization.
@@
This leads us to the following discretized relaxation of the problem. Let $t_{i}\to 0$ and set $\mu_{0}=N(x^{*},\alpha I)$, an isotropic initialization. If we compute the initial $\hat{\mathcal{L}}(\mu_{0})$, we may hope that, as long as $\mu_{1}$ represents a sufficiently comparable distribution, $\hat{\mathcal{L}}(\mu_{0})\approx\hat{\mathcal{L}}(\mu_{1})$. Thus, we can compute
$\Sigma_{1}=\operatorname{argmax}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{0})\Sigma]\leq t_{1}}\log\det\Sigma=t_{1}\hat{\mathcal{L}}(\mu_{0})^{-1}/D.$
@@
Let $t_{i}\to 0$, fix $\Sigma_{0}$, $\mu_{i}=N(x^{*},\Sigma_{i})$, $\hat{\mathcal{L}}_{i}:=\hat{\mathcal{L}}(\mu_{i}):=\int\hat{\nabla}f\hat{\nabla}f^{T}\,d\mu_{i}$, and construct $\hat{\nabla}f=\hat{P}_{\Sigma_{i}^{-1/2}}(f)$. Define the recursive kernel learning algorithm $\Sigma_{i+1}=t_{i+1}\hat{\mathcal{L}}_{i}^{-1}/D$. Then, equivalently,
$\Sigma_{i+1}$ $=\operatorname{argmax}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{i})\Sigma]\leq t_{i+1},\Sigma\succeq 0}\log\det\Sigma$ (6)
$=\operatorname{argmin}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{i})\Sigma]=t_{i+1},\Sigma\succeq 0}\operatorname{tr}[\hat{\mathcal{L}}(\mu_{i})\Sigma]+\frac{1}{n\sqrt{\det\Sigma}}.$
We note that in Algorithm 1, we use local linear regression to better fit into the EGOP estimation loop. We expect that a suitable extension of our theory is possible for higher-order polynomials, allowing for more complex local regressions in Algorithm 1. See Section 8 for further details.
From this framework we have developed, it is clear that there are many possible schemes to optimize the bias–variance trade-off, and we discuss this in more detail in Section 8. We choose to focus on this alternating optimization scheme because it strongly resembles algorithms present in the existing literature, and is simple to implement.
6 Guarantees
In this section, we demonstrate concrete theoretical guarantees for estimation via Local EGOP Learning. In our theory, we do not consider the problem of EGOP estimation, and instead consider an iterative scheme in which we can query an oracle for the EGOP associated to a particular choice of covariance. For our experiments, we compute empirical AGOPs fit via the proposed iterative estimation procedure, leading to sharp estimation quality. We leave it to future work to close this theoretical gap, although we discuss it in greater depth in Appendix C. See Appendix B for proofs of all following arguments.
6.1 Taylor Expansion
Up to this point, our prevailing perspective is that the target distributions $\mu$ matter up to their first and second moments, passing the optimization problem from the space of measures to the PSD cone, and informing a Gaussian relaxation. We close the loop to EGOP estimation by additionally Taylor expanding $\mathcal{L}(\mu)$ up to second order, reducing the proposed iteration scheme entirely to a recurrence in $\Sigma$.
@@ 6.1 Taylor Expansion
$\Sigma_{i+1}$ $=t_{i+1}(\beta\mathcal{L}(\mu_{i})+(1-\beta)\mathcal{L}(\mu_{i-1}))^{-1}$ (7)
$=t_{i+1}(gg^{T}+\beta[H\Sigma_{i}H+T_{g}(\Sigma_{i})+R(\Sigma_{i})],$
$\quad\quad+(1-\beta)[H\Sigma_{i-1}H+T_{g}(\Sigma_{i-1})+R(\Sigma_{i-1})])^{-1}.$
We initialize this recurrence with the isotropic choice $\Sigma_{0}=\alpha I$, $\Sigma_{1}=t_{1}\mathcal{L}(\mu_{0})^{-1}$, $\Sigma_{2}=t_{1}\mathcal{L}(\mu_{1})^{-1}$, then apply momentum following this step.
6.2 Full Rank Hessian
We begin our analysis with the generic setting, where $H$ is full-rank. In this setting, the Taylor expansion of $\mathcal{L}(\mu)$ is dominated by a rank 1 matrix $gg^{T}$ and a linear term $H\Sigma H$, which combine to yield a full rank matrix. We show that Local EGOP Learning results in a marked efficiency gain over generic isotropic kernel methods. Indeed, we argue that $g^{T}\Sigma_{i}g=\Theta(t_{i})$, while for $v\perp g,$ $v^{T}\Sigma_{i}v=\Theta(\sqrt{t_{i}})$, thus achieving a second-order anisotropy. The noisy manifold setting implies a rank deficiency of $H$, and we will show in the next section how this leads to further efficiency gains.
@@ 6.2 Full Rank Hessian
Fix a schedule for $t_{i}\to 0$ that is no faster than geometric, so that $t_{i+1}/t_{i}$ is larger than a fixed threshold. We investigate a considerable reduction of the problem, the one dimensional recurrences
$a_{i+1}$ $=\frac{t_{i+1}}{\beta a_{i}+(1-\beta)a_{i-1}},$ (8)
$b_{i+1}$ $=\frac{t_{i+1}}{c+\beta b_{i}+(1-\beta)b_{i-1}}.$ (9)
It is easy to see that $b_{i}=O(t_{i})$, as if $b_{i}$ decays at all, then we have $b_{i}=t_{i}/c+o(t_{i})$. One can also readily observe, at least informally, the rate of decay of $a_{i}$. Assume that $a_{i}=\Theta(t_{i}^{\zeta})$. By the recurrence relation, we then have $a_{i}=\Theta(t_{i}^{1-\zeta})$, hence these coincide at the second order anisotropy $\alpha=1/2$. As our matrix recurrence is primarily given by a homogeneous part $(H\Sigma H)$ plus a rank deficient constant part $(gg^{T})$, we argue in Appendix B that the $b_{i}$ correspond well to the $g\times g$ block of $\Sigma_{i}$, and $a_{i}$ to $g^{\perp}\times g^{\perp}$.
@@
We now take a moment to highlight the necessity of the momentum term $\beta>0$. Consider, again, the sequence $a_{i}$ from Example 1, and suppose we set $\beta=1$. For convenience, let us assume $\sqrt{t_{i+1}}=\sqrt{t_{i}}r$, an exact geometric schedule. We reparameterize, writing
$a_{i+1}=\frac{t_{i+1}}{a_{i}}\Longleftrightarrow a^{\prime}_{i+1}=\frac{1}{a^{\prime}_{i}}$
for $a_{i}^{\prime}:=\frac{a_{i}}{\sqrt{rt_{i}}}$. Clearly, $a_{i}^{\prime}$ has no limit, as it simply oscillates wildly about $1$. The remainder terms, as well as the estimation error, make the $\sqrt{t_{i}}$ rate become unstable, leading to potential catastrophic failure. We demonstrate the resulting localizations with and without momentum in Figure 4.
Figure 4: Comparison of the eigenvalue decay of $\Sigma_{i}$ for Local EGOP Learning with and without momentum. On the left, we set $\beta=0.7$, on the right $\beta=1$ (no momentum). Without momentum, the second order eigenvalues are prone to wild oscillations.
@@
Let $H$ be invertible and $\Sigma_{i}$ satisfy the recurrence in equation 7, $\sqrt{t_{i}}=\alpha r^{i}$, $0<r<1,\ \beta>0$. Then, for $\alpha$ sufficiently small, $g^{T}\Sigma_{i}g=\Theta(t_{i})$, and for $v\perp g$, $v^{T}\Sigma_{i}v=\Theta(\sqrt{t_{i}})$. It follows that, selecting $i_{n}=\frac{4\log n}{(D+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{\Sigma_{i_{n}}}(f)-f)^{2}\,d\mu_{i_{n}}\right]=O\left(n^{-\frac{4}{D+5}}\right).$
@@
6.3 Noisy Manifold
In the noisy manifold setting we strengthen Lemma 4.

### v3

@@ 4.2 Bias Control via EGOP
In this section, we develop a theoretical framework for adaptive kernel learning. Define the empirical kernel estimator for predicting the value of $f$ at $x^{*}$ by
$\hat{P}_{M_{t}}(f)$ $:=\frac{1}{\hat{C}_{M_{t}}}\frac{1}{n}\sum_{i=1}^{n}k(\|M_{t}^{1/2}[X_{i}-x^{*}]\|/\sqrt{2h_{t+1}})Y_{i},$
$\hat{C}_{M_{t}}$ $:=\frac{1}{n}\sum_{i=1}^{n}k(\|M_{t}^{1/2}[X_{i}-x^{*}]\|/\sqrt{2h_{t+1}}).$
Its continuous counterpart is
$P_{M_{t}}(f)$ $:=\frac{1}{C_{M_{t}}}\int k(\|M_{t}^{1/2}[y-x^{*}]\|/\sqrt{2h_{t+1}})f(y)\,dP(y),$
$C_{M_{t}}$ $:=\int k(\|M_{t}^{1/2}[y-x^{*}]\|/\sqrt{2h_{t+1}})\,dP(y).$
@@
The bandwidth $h$ is suppressed in our notation as it will be pre-set at each iteration of our algorithm. To assess the quality of this predictor, we can measure its bias $\int(P_{M_{t}}(f)-f)^{2}\,d\mu$ on a target distribution $\mu$. When $M$ and $\mu$ are chosen in tandem, the bias can be related to a Dirichlet form in terms of the EGOP. To make this explicit, we introduce the EGOP form
@@
Let $M_{t}\succeq 0$, and set $\mu_{t+1}=N(0,h_{t+1}M_{t}^{-1})$. Then,
$\int(P_{M_{t}}(f)-f)^{2}\,d\mu_{t+1}=O(W(\mu_{t+1})).$
In the anisotropic metric $M_{t}$, the variation of $f$ is given by $\mathcal{L}(\mu_{t+1})$, and the variance of $x$ is given by $\Sigma(\mu_{t+1})$. If these matrices are close to low-rank and perpendicular, then the bias of the kernel estimator is controlled by $W(\mu_{t+1})=\langle\mathcal{L}(\mu_{t+1}),\Sigma(\mu_{t+1})\rangle_{F}$, and will be nearly 0. Of critical importance are the first and second moments of the target distribution $\mu$, motivating our Gaussian convention, although Gaussianity is not essential to achieve the desired bound. We note that this general structure, of metrizing by the inverse covariance, appears in disparate literatures, from classical local feature estimators (Schmid and Mohr, 1997), to VAEs (Chadebec and Allassonnière, 2022), and Langevin preconditioning (Cui et al., 2024).
@@
Let $M_{t}\succeq 0$, $\Sigma_{t+1}=h_{t+1}M_{t}^{-1}$. Then $\mathbb{E}\left[(\hat{P}_{M_{t}}(f)-P_{M_{t}}(f))^{2}\right]=O\left(1/(n\sqrt{\det\Sigma_{t+1}})\right).$
@@
Let $M_{t}\succeq 0$, and set $\Sigma_{t+1}=h_{t+1}M_{t}^{-1}$, $\mu_{t+1}=N(x^{*},\Sigma_{t+1})$. Then,
$\mathbb{E}\left[\int(\hat{P}_{M_{t}}(f)-f)^{2}\,d\mu_{t+1}\right]=O\left(\frac{1}{n\sqrt{\det\Sigma_{t+1}}}+W(\mu_{t+1})\right).$
In particular, if $W(\mu_{t})=O(h_{t})$ and $\det\Sigma_{t}=O(h_{t}^{q})$ then the asymptotic learning rate is
$\mathbb{E}\left[\int(\hat{P}_{M_{t}}(f)-f)^{2}\,d\mu_{t+1}\right]=O\left(n^{-\frac{1}{1+q/2}}\right).$
4.3 Recursive Kernel Learning as Variance Minimization
@@
We will now show how recursive kernel learning, combined with our basic kernel bounds, constitutes a greedy variance reduction strategy. Intuitively, the goal is to induce anisotropy in the kernel, stretching or steering it to include additional low-bias points that would otherwise be discarded with an isotropic metric. The inclusion of these additional high-quality points is exactly a reduction of variance, as including more data results in tighter concentration to the mean for a local averaging estimator.
@@
We now make precise how one can arrive at this by leveraging Theorem 1. Ideally, as spelled out in this result, we would like to select $\mu_{t}$ such that $W(\mu_{t})\to 0$, and that $\det\Sigma(\mu_{t})$ is as large as possible. In other words, if we specify the rate $W(\mu_{t})=\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]=h_{t}\to 0$, it would be optimal to select the largest possible region that achieves this bias, meaning we would like to select
$\mu_{t}=\operatorname{argmax}_{\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]\leq h_{t}}\log\det\Sigma(\mu_{t}).$
@@
This however cannot be solved, as we lack precise knowledge of the target function $f$. Moreover, $\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]\leq h_{t}$ presents a challenging feasible set, as it is posed over a space of probability distributions, thus making it non-Euclidean. Our Gaussian ansatz reduces the problem to optimizing a finite parameterization, though the optimization is non-convex.
@@
This leads us to the following discretized relaxation of the problem. Set $\mu_{0}=N(x^{*},\alpha I)$, an isotropic initialization. If we compute the initial $\hat{\mathcal{L}}(\mu_{0})$, we may hope that, as long as $\mu_{1}$ represents a sufficiently comparable distribution, $\hat{\mathcal{L}}(\mu_{0})\approx\hat{\mathcal{L}}(\mu_{1})$. Thus, we can compute
$\Sigma_{1}=\operatorname{argmax}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{0})\Sigma]\leq t_{1}}\log\det\Sigma=h_{1}\hat{\mathcal{L}}(\mu_{0})^{-1}/D.$
@@
Let $h_{t}\to 0$, fix $\Sigma_{0}$, $\mu_{t}=N(x^{*},\Sigma_{t})$, $\hat{\mathcal{L}}_{t}:=\hat{\mathcal{L}}(\mu_{t})$, and construct $\hat{f}=\hat{P}_{\Sigma_{t}^{-1/2}}(f)$. Let $\Sigma_{t}$ satisfy the recursive kernel recurrence $\Sigma_{t+1}=h_{t+1}\hat{\mathcal{L}}_{t}^{-1}/D$. Then, equivalently, optimizing over $\Sigma\succeq 0$,
$\Sigma_{t+1}$ $=\operatorname{argmax}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{t})\Sigma]\leq h_{t+1}}\log\det\Sigma$ (4)
$=\operatorname{argmin}_{\operatorname{tr}[\hat{\mathcal{L}}(\mu_{t})\Sigma]=h_{t+1}}\operatorname{tr}(\hat{\mathcal{L}}(\mu_{t})\Sigma)+\frac{1}{n\sqrt{\det\Sigma}}.$
We note that in Algorithm 1, we use local linear regression to better fit into the EGOP estimation loop, and we develop corresponding theory in Appendix C.
From this framework we have developed, it is clear that there are many possible schemes to optimize the bias-variance trade-off, and we discuss this in more detail in Section 6. We choose to focus on this alternating optimization because it strongly resembles algorithms present in the existing literature, and is simple to implement.
4.4 Bias-Variance trade-off – Convergence at intrinsic rate
In this section, we demonstrate concrete theoretical guarantees for estimation via Local EGOP Learning. For ease of presentation, in the main text we consider an iterative scheme in which an oracle EGOP is queried for a particular choice of covariance. In Appendix C we develop theory towards EGOP approximation, showing that anisotropic smoothing is compatible with recursive kernel learning under standard nonparametric estimation rates. For our experiments, we compute empirical AGOPs fit via the proposed iterative estimation procedure, leading to sharp estimation. See Appendix B for proofs of all following arguments.
4.4.1 Taylor Expansion
Up to this point, our prevailing perspective is that the target distributions $\mu$ matter up to their first and second moments, passing the optimization problem from the space of measures to the PSD cone, and facilitating a Gaussian relaxation. We close the loop to EGOP estimation by additionally Taylor expanding $\mathcal{L}(\mu)$ up to second order, reducing the proposed iteration scheme entirely to a recurrence in $\Sigma$.
@@ 4.4.1 Taylor Expansion
$\Sigma_{t+1}$ $=h_{t+1}(\beta\mathcal{L}(\mu_{t})+(1-\beta)\mathcal{L}(\mu_{t-1}))^{-1}$
$=h_{t+1}(gg^{T}+\beta[H\Sigma_{t}H+T_{g}(\Sigma_{t})+R(\Sigma_{t})],$
$\quad\quad+(1-\beta)[H\Sigma_{t-1}H+T_{g}(\Sigma_{t-1})+R(\Sigma_{t-1})])^{-1}.$
We initialize this recurrence with the isotropic choice $\Sigma_{0}=\alpha I$, $\Sigma_{1}=t_{1}\mathcal{L}(\mu_{0})^{-1}$, $\Sigma_{2}=t_{1}\mathcal{L}(\mu_{1})^{-1}$, then apply the momentum $\beta$ following this step.
4.4.2 Full Rank Hessian
We begin our analysis with the generic setting, where $H$ is full-rank. Here, the Taylor expansion of $\mathcal{L}(\mu)$ is dominated by a rank 1 matrix $gg^{T}$ and a linear term $H\Sigma H$, which combine to yield a full rank matrix. We argue that $g^{T}\Sigma_{t}g=\Theta(h_{t})$, while for $v\perp g,$ $v^{T}\Sigma_{t}v=\Theta(\sqrt{h_{t}})$, thus achieving a second-order anisotropy.
@@ 4.4.2 Full Rank Hessian
Fix a schedule for $h_{t}\to 0$ such that $h_{t+1}/h_{t}$ is larger than a fixed threshold. Consider the one dimensional recurrences
$a_{t+1}$ $=\frac{h_{t+1}}{\beta a_{t}+(1-\beta)a_{t-1}},$ (5)
$b_{t+1}$ $=\frac{h_{t+1}}{c+\beta b_{t}+(1-\beta)b_{t-1}}.$ (6)
It is easy to see that $b_{t}=O(h_{t})$, as if $b_{t}$ decays at all, then we have $b_{t}=h_{t+1}/c+o(h_{t})$. Assume that $a_{t}=\Theta(h_{t}^{\zeta})$. By the recurrence relation, we then have $a_{i}=\Theta(h_{t}^{1-\zeta})$, and so these coincide at the second order anisotropy $\zeta=1/2$. As our matrix recurrence is primarily given by a homogeneous part $(H\Sigma H)$ plus a rank deficient constant part $(gg^{T})$, we argue in Appendix B that the $b_{t}$ correspond well to the $g\times g$ block of $\Sigma_{t}$, and $a_{i}$ to $g^{\perp}\times g^{\perp}$.
@@
We now take a moment to highlight the necessity of the momentum term $\beta\in(0,1)$. Consider, again, the sequence $a_{t}$ from Example 1, and suppose we set $\beta=1$. For convenience, let us assume $\sqrt{h_{t+1}}=\sqrt{h_{t}}r$, an exact geometric schedule. We reparameterize, writing
$a_{t+1}=\frac{h_{t+1}}{a_{t}}\Longleftrightarrow a^{\prime}_{t+1}=\frac{1}{a^{\prime}_{t}}$
for $a_{i}^{\prime}:=\frac{a_{i}}{\sqrt{rt_{i}}}$. Clearly, $a_{t}^{\prime}$ has no limit, as it oscillates about $1$. In the corresponding matrix recurrence, the remainder terms, as well as the estimation error, make the $\sqrt{h_{t}}$ rate become unstable. We demonstrate the resulting localizations with and without momentum in Figure 4.
Figure 4: Comparison of the eigenvalue decay of $\Sigma_{t}$ for Local EGOP Learning with and without momentum. On the left, we set $\beta=0.7$, on the right $\beta=1$ (no momentum). Without momentum, the second order eigenvalues are prone to wild oscillations.
@@
Let $H$ be invertible and $\Sigma_{t}$ satisfy the recurrence in equation 2, $\sqrt{h_{t}}=\alpha r^{t}$, $0<r<1,\ \beta>0$. Then, for $\alpha$ sufficiently small, $g^{T}\Sigma_{t}g=\Theta(h_{t})$, and for $v\perp g$, $v^{T}\Sigma_{t}v=\Theta(\sqrt{h_{t}})$. It follows that, selecting $t_{n}=\frac{4\log n}{(D+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{M_{t_{n}}}(f)-f)^{2}\,d\mu_{t_{n}}\right]=O\left(n^{-\frac{4}{D+5}}\right).$
@@
4.4.3 Noisy Manifold
In the noisy manifold setting, we strengthen Lemma 3.